5 years ago, Jay was seven years older than Mary. In five years, Mary will be half as old as Jay (or Jay will be twice as old as Mary). How old is Mary now?
Answer by bluemockingjay7(41)

present age:
Mary = x
Jay = 2x
age 5 years ago:
Mary = x - 5
Jay = 2x - 5
Equation:
2x - 5 = 7(x - 5)
2x - 5 = 7x - 35
2x - 7x = -35 + 5
-5x = -30
-5x/5 = -30/-5
x = 7
present age:
Mary = 7
Jay = 2(7) = 14
age 5 years ago:
Mary = 7 - 5 = 2
Jay 14 - 5 = 9